# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v1

@@ 3.1 Behavior can be different in naturalistic experiments
The starting point for many cognitive analyses is behaviors on a task designed to isolate some cognitive capability. However, there are many cases in which task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted task setting unless we also consider a broading range of variation. In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
@@
Example 1: moral behavior can change when moving from textual vignettes to virtual reality. Recent research has shown that behavioral choices can change when moving from text to virtual reality as it can mark a shift from moral judgment (i.e. deciding whether things are moral) to moral action (Francis et al.,, 2017). When presented with the classic trolley dilemma in text form, only 40% of participants say they would pull a switch to divert a trolley that would kill one person to save five others. However, when the same scenario is presented in virtual reality where participants must simulate the action, 55% choose to pull the switch. This difference becomes even more pronounced in the “footbridge” variant of the dilemma. In the text version, only 10% of participants say they would push someone off a bridge to save five others. Yet when able to simulate this action in virtual reality, 65% of participants choose to push the person. These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
@@
Example 2: memory performance differs between naturalistic search and explicit memorization. When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020). This effect extends to spatial memory as well. Not only do participants better remember the identity of objects encountered during search, they also show more accurate memory for object locations—placing objects closer to their original positions when they were initially seen during search versus explicit memorization. These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory naturally operates during everyday behavior. Additionally, it suggests that computational models of memory may need to be revised to account for these differences in naturalistic settings. This is supported by research showing that people display detailed memory of object and spatial information in real-world scenes (Bainbridge et al.,, 2019).
@@
Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice tasks. Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021). When mice are tested in traditional two-alternative forced choice (2AFC) tasks, they typically require 10,000 trials over 3-6 weeks of training to reach asymptotic performance of only about 67% accuracy. However, when allowed to freely explore a labyrinth environment (similar to a natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021). In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, discovering and optimizing their path to a water reward after only about 10 reward experiences. The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in traditional 2AFC paradigms. This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that natural learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
@@ 3.2 Neural systems can operate differently under naturalistic conditions
In addition to behavioral differences like those reviewed above, naturalistic conditions can result in changes to neural coding and computation. Thus, in order to arrive at a complete understanding of neural function, we need to consider the responses of the system across a range of simpler and more naturalistic settings. In this section, we highlight some examples of different computations or computational roles that arise when moving from standard paradigms to more naturalistic ones.
@@
Example 1: the model of superior colliculus shifted from a model of eye movement to one of integrating sensory inputs for body control. Cisek, (1999) suggests that the focus on computational behavior as an input-output mapping neglects the fundamental fact that natural intelligence evolved not to produce single responses, but for closed-loop control in an environment. This setting yields a rather different interpretation of the system’s representations and processes. More generally, many researchers have argued that cognition cannot be understood completely outside its embodiment and the environment in which cognitive processes are instantiated (e.g., Newen et al.,, 2018). For example, Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free to body-free settings when studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Cisek and Green,, 2024). That is, as we increased the naturalism of the experimental conditions that we used to study monkeys, we arrived at a more complete model of superior colliculus.
@@
Example 2: visual processing systems operate differently during natural viewing compared to passive viewing paradigms. Traditional research has shown that one of the earliest visual responses in the brain (known as P100/M100) occurs about 100ms after a person’s eyes land and fixate on a new location (Vinje and Gallant,, 2000). However, this finding comes primarily from experiments where participants passively view images while keeping their eyes still. In contrast, natural vision involves actively moving our eyes multiple times per second to sample information from different parts of a scene. When the researchers examined brain responses during this more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) starts, not when it ends at the new fixation point (Amme et al.,, 2024). Furthermore, they discovered that the neural patterns during active viewing were opposite (anti-correlated) to those seen during passive viewing, suggesting fundamentally different processing mechanisms. These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing actually operates during natural behavior.
@@
Example 3: model-based learning systems are used for learning moral judgments, but model-free learning are used for learning to avoid harming others. Despite moral reasoning being strongly associated with model-based learning systems in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020). Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013). However, when participants engage in tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning. Similar to our moral judgment vs. moral action example from §3.1, this shows that when an experiment becomes more naturalistic (not just reasoning about morality but learning moral behavior), neural systems can operate in unexpected ways.
@@
Finally, naturalistic paradigms can likewise introduce challenges that are important to understanding functioning of the system as a whole. In this section, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can be useful for disentangling underlying mechanisms, and is important for achieving generalizable understanding.
@@
Example 1: Different learning mechanisms only yield distinct predictions under working memory load. Overly-simplified tasks can yield aliasing of different solutions, where many different computational approaches yield the same behavior. Collins, (2024) presents an example, by showing that what seems superficially to be implemented as a standard reinforcement learning algorithm, on closer examination may instead be implemented through a combination of working memory and simpler value-free associative learning. [^1] Whether this alternative counts as an alternative implementation of RL or not is orthogonal to our argument; the important point is that the two models yield distinct predictions in some regimes but not others. However, the difference between these two implementations cannot be identified in standard task settings, as Collins notes: “Even simple tasks designed to elicit a target process (such as bandit tasks for RL) recruit multiple other processes, but those processes may be unidentifiable in such tasks. Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the stimulus set size within learning blocks to a handful of objects, rather than restricting to two or three—i.e., more broadly covering the range of set sizes more clearly disentangles the learning processes. Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them. By contrast, if we impose a broad range of evaluations on the system (cf. Nau et al.,, 2024)—and especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence, as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
@@
Example 2: fear conditioning. Traditional fear conditioning research has provided fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014). Through carefully controlled laboratory studies involving paired stimuli and outcomes, we have developed a rich mechanistic understanding of basic fear learning processes. These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events like combat exposure (Engelhard et al.,, 2008). However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes. For example, a student may develop fear of exams after receiving a single poor grade weeks after taking the test, without requiring repeated negative experiences (Mobbs et al.,, 2021). This type of naturalistic learning involves delayed delayed and sparse feedback for choices, i.e. it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018). Credit assignment describes how an agent credits choices (e.g. studying, socializing, diet, etc.) as responsible for future events (e.g. a poor grade). This is not to say humans have optimal credit assignment but that they have some mechanism for it and we are empowered to study it with appropriately naturalistic experimental settings. Clinical researchers have additionally pointed out that increased ecological validity can increase the applicability of our experimental results as the gap between lab-based fear conditioning and real-world fear has been a barrier to successful treatment of pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
@@
Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models. There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that is different from the rest of the objects in a set) recruit the Medial Temporal Lobe (MTL). Specifically, some researchers (e.g Murray et al.,, 2007) have suggested that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (and thus MTL lesions impair performance specifically on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009). Recently, stimulus-computable models have shed new light on this debate. Specifically, Bonnen et al., (2021); Bonnen and Eldridge, (2023) used pretrained convolutional vision models to directly process all the raw stimuli used in prior experiments, and evaluated whether the “oddity” stimulus could be distinguished from the rest of the set in the representation space of these models. By doing so, Bonnen et al. showed that some of the purportedly “complex” stimuli from some of the prior experiments could nevertheless be distinguished with the vision model alone, while the complex stimuli from other experiments could not. Furthermore, the purportedly “complex” stimuli that in fact could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL. Thus, the way that prior works operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings. This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms. This naturalistic approach can elucidate important differences among experimental paradigms, and clarify conceptual understanding.
@@ 4 The (surprising) benefits of learning with naturalistic data
The standard lesson in science is that we must simplify an experimental setting so that we can better arrive at a causal conclusion. However, we have already seen examples illustrating that more is different (Anderson,, 1972)—certain properties of intelligence may only emerge in more “complicated” naturalistic settings. In this section, we detail examples where naturalistic data itself seems to play an importance role in producing the empirical phenomena of intelligence that we wish to study.
Many of the findings we will discuss here will be modern results from AI. One common trend is that there is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively better results than learning in simpler settings. While we do not yet know the implications for natural intelligence, they open interesting questions and challenge prior assumptions. In particular, we argue that cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data. (Note that while we draw many case studies from particular areas of deep learning, we believe that some of these factors could likewise contribute to generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
@@ 4.1 Learning enables models that can operate on natural data
Perhaps the first surprising benefit of (deep) learning is that it enables a system to operate over natural data. Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models. However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply by being trained to make predictions about this data. Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014). Here, we review some of this work and how learning-based methods can enable more flexible cognitive models that operate on natural data.
@@
Learning-based vision models can operate on natural data and predict neural data (surprisingly) well. There is a rich history of learning-based methods revealing how neural networks can develop internal representations from pixel-based images that strongly correspond to neural data. A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional neural network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder). Prior to this, V1 activations were modeled using analytically defined Gabor filters. This breakthrough demonstrated that V1-like receptive fields could arise naturally from optimization to environmental statistics of edges at varying frequencies and orientations. Building on this foundation, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks. A significant advance came when Yamins and DiCarlo, (2016) showed that task-optimized deep convolutional neural networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models. Recent research has revealed an even more striking finding: these task-optimized networks represent images in a latent space whose geometry has remarkable correspondence to brain topography (Doshi and Konkle,, 2023). When fitting a 2D sheet to image representations such that neighboring points correspond to nearby points in the original high-dimensional space, the resulting sheet shows striking similarities to actual brain topography. Thus, learning-based approaches, particularly neural networks, not only enable effective visual processing but also provide surprisingly accurate predictions of brain organization and function.
@@
Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning. Recent work has shown how large language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning (Ellis,, 2024). Hand-designed symbolic hypothesis spaces don’t always transfer across domains.  Ellis, (2024) showed that one can circumvent these challenges by leveraging large language models to enable use of natural language as a flexible hypothesis space with a data-driven prior tuned to human judgments. By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains. This overcomes traditional limitations of pure symbolic program synthesis approaches which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004), where when deciding the relevant variables for a problem, we might accidentally choose a framing that doesn’t generalize to other settings of interest. The integration of learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modellers.

### v2

@@ 3.1 Behavior can be different in naturalistic experiments
Many cognitive analyses begin by studying a task designed to isolate some cognitive capability. However, in many cases task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted setting unless we also consider a broader range of variation. In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
@@
Example 1: moral behavior can change when moving from textual vignettes to virtual reality. Recent research shows that behavioral choices can change when moving from text to virtual reality, by driving a shift from moral judgment (i.e. deciding whether things are moral) to moral action (Francis et al.,, 2017, see Figure 4 A). When participants were presented the classic trolley dilemma in text form, only 40% said they would pull a switch to divert the trolley, thus killing one person to save five others. However, when the same scenario was presented in virtual reality, where participants must simulate the action, 55% chose to pull the switch. The difference was even more pronounced in the “footbridge” variant of the dilemma. In the text version, only 10% of participants said they would push someone off a bridge to save five others. However, in virtual reality, 65% of participants choose to push the person. These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
@@
Example 2: memory performance differs between naturalistic search and explicit memorization. When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020). This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition. These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior. Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
@@
Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks. Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021). When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy. However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021). In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences. The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in 2AFC paradigms. This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
@@ 3.2 Neural systems can operate differently under naturalistic conditions
In addition to behavioral differences like those reviewed above, naturalistic conditions can alter neural coding and computation. In this section, we highlight some examples where different computations or roles arise when moving from standard paradigms to more naturalistic ones.
@@
Example 1: the role of superior colliculus in eye movements vs. integrating sensory inputs for body control. Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B). That is, as we increased the naturalism of the experimental conditions, we arrived at a more complete model of superior colliculus. This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek,, 1999). This perspectives yields a different interpretation of the system’s representations and processes. More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al.,, 2018).
@@
Example 2: visual processing differs during natural viewing compared to passive viewing paradigms. One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant,, 2000). However, this finding comes primarily from experiments where participants passively view images while fixating. In contrast, natural vision involves frequent active eye movements to sample information throughout a scene. When Amme et al., (2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point. Furthermore, the neural patterns during active viewing were anti-correlated to those seen during passive viewing, suggesting fundamentally different processing mechanisms. These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing operates during natural behavior.
@@
Example 3: model-based systems are used for learning moral judgments, but model-free systems are used for learning to avoid harming others. Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020). Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013). However, when participants perform tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning. As in §3.1, this illustrates that increasing naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
@@
Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable understanding.
@@
Example 1: Different learning mechanisms only yield distinct predictions under working memory load. One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior. Collins, (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory. [^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes. They note that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes. Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them. By contrast, by imposing a broad range of evaluations on the system (cf. Nau et al.,, 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence (more constraints on the mapping make it less ambiguous), as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
@@
Example 2: fear conditioning. Traditional fear conditioning research offers fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014). Through carefully controlled laboratory studies, we have developed a rich mechanistic understanding of basic fear learning processes. These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events (Engelhard et al.,, 2008). However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes (Figure 4 C). For example, a student may develop a fear of exams after receiving a single poor grade weeks after taking a test, without requiring repeated negative experiences (Mobbs et al.,, 2021). This naturalistic learning involves delayed and sparse feedback. Thus, it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018)—a mechanism for crediting choices (e.g. studying or sleep) as responsible for future events (e.g. a poor grade). Naturalistic settings enable studying important functions such as mechanisms for credit assignment. Increased ecological validity can thereby increase the applicability of our findings; indeed, the gap between lab-based fear conditioning and real-world fear has been a barrier to treating pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
@@
Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models. There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL). Some researchers (e.g Murray et al.,, 2007) suggest that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (thus MTL lesions selectively impair performance on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009). Recently, stimulus-computable models have shed new light on this debate. Specifically, Bonnen et al., (2021); Bonnen and Eldridge, (2023) used pretrained convolutional vision models to directly process all the raw stimuli used in prior experiments, and evaluated whether the “oddity” stimulus could be distinguished from the rest of the set in the representation space of these models. By doing so, Bonnen et al. showed that purportedly “complex” stimuli that could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL. This shows that the way that prior works had operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings. This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms. This naturalistic approach can elucidate important differences among experimental paradigms, and clarify conceptual understanding.
@@ 4 The (surprising) benefits of learning with naturalistic data
A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion. However, we have reviewed examples illustrating that more is different (Anderson,, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings. Here, we detail examples where naturalistic data itself plays an important role in producing the empirical phenomena of intelligence.
Many results we discuss are drawn from modern AI research. One common trend is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively different generalization than learning in simpler settings. While we do not yet know the full implications for natural intelligence, these results open interesting questions and challenge prior assumptions. In particular, cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data. (While we draw case studies from deep learning, we believe that similar factors could likewise affect generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
@@ 4.1 Learning enables models that can operate on natural data
One surprising benefit of learning is that it can enable a system to operate over natural data. Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models. However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply through learning to make predictions about those data. Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014). Here, we review some of this work and how learning-based methods can enable more flexible cognitive models that operate on natural data.
@@
Learning-based vision models can operate on natural data and predict neural data (surprisingly) well. A rich history of work reveals how neural networks can learn internal representations from natural images that strongly correspond to neural data. A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder). This breakthrough demonstrated that V1-like receptive fields could arise from optimization for natural environmental statistics. Subsequently, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks. Later, Yamins and DiCarlo, (2016) showed that deep convolutional networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models. Strikingly, task-optimized networks trained to self-organize represent images in a latent space whose geometry corresponds closely to brain topography (Doshi and Konkle,, 2023, Figure 5 A). Thus, learning-based approaches not only enable effective visual processing but also can provide surprisingly accurate predictions of brain organization and function.
@@
Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning. Recently, Ellis, (2024) showed how language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning. Language models enable using natural language as a flexible hypothesis space with a prior derived from human data. By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains. This overcomes traditional limitations of hand-designed symbolic hypothesis spaces and pure program synthesis approaches, which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004) in hypothesis space design.Integrating learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modelers.